n 1980, major US banks offered interest rates as high as 15% per year on savings accounts offered to consumers. By 2000, that rate had dropped to 5% per year. On the other hand, in 1980 the US inflation rate was about 14% per year, whereas in 2000, the US inflation rate was only 3% per year. a. What was the real interest rate in 1980? What was the real interest rate in 2000?

a) What was the real interest rate in 1980? What was the real interest rate in 2000?

b) What was the nominal interest rate in 1980? What was the nominal interest rate in 2000?

c) For a consumer, which year – 1980 or 2000 – was a better time to put money in the bank?


1
Expert's answer
2021-04-06T07:02:35-0400

a) The real interest rate in 1980 was r = 15 - 14 = 1%.

The real interest rate in 2000 was r = 5 - 3 = 2%.

b) The nominal interest rate in 1980 was 15%.

The nominal interest rate in 2000 was 5%.

c) For a consumer, year 2000 – was a better time to put money in the bank, because the real interest rate was higher.